OpenStudy (anonymous):

A rectangle is drawn so that the width is 5 feet shorter than the length. The area of the rectangle is 50 square feet. Find the length of the rectangle. feet

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

let length = l width = w therefore according to the question we have w= l-5 Area of rectangle = l*w i.e. 50 = l(l-5) i.e. \[\huge 50 = l^2 - 5l \rightarrow l^2- 5l -50 =0\] \[\huge \rightarrow l^2- 10l +5l -50 =0\] \[\huge \rightarrow l(l- 10) +5(l -10) =0\] \[\huge \rightarrow l(l- 10) (l+5) =0\] i.e. either l-10=0 or l+5 =0 i.e. l= 10 or \[\huge l =-5\] we will take the positive value of l hence length l of the rectangle is 10 feet

OpenStudy (anonymous):

@cancherolugano

OpenStudy (anonymous):

Sorry, remove l from the 7th line from down outside of the bracket

OpenStudy (anonymous):

No problem thank you as always